OpenStudy (anonymous):

2 men start from 2 places 600km apart and travel towards each other, the first travelling 5km/h faster. they meet after 8 hours.find te speed of the fastest man

OpenStudy (anonymous):

600/8 = 75 since they are moving in opposite direction their relative velocity is equal to sum of individual velocity

OpenStudy (anonymous):

let their speeds be x and x-5 units respectively then by the prob x+x-5=600/8 2x-5 =75 2x=80 x=40 so speed of the fastest one is 40km/h

OpenStudy (anonymous):

thank you matricked

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

how about this one. mary noticed that she had 104 coins made up entirely of 20c and 5c coins that totalled exactly $10. how many of each type of coin did she have?

OpenStudy (wolf1728):

20¢ coins??? 19th century USA isn't it?

OpenStudy (anonymous):

Aussie coins

OpenStudy (wolf1728):

oops sorry for my USA centric outlook :-(

OpenStudy (aum):

Let number of 20 cent coins be x and the number of 5 cent coin be y x + y = 104 x/5 + y/20 = 10 Solve for x and y.

OpenStudy (aum):

Or if you would like you can write the second equation as 0.2x + 0.05y = 10

OpenStudy (anonymous):

Thanks! One last question. John is twice as old as Mary. five years ago he was three times as old. What is Mary's age now?

OpenStudy (aum):

Let M be Mary's age and J be John's age J = 2M 5 years ago, John would have been (J-5) years old and Mary (M-5) years old J - 5 = 3(M - 5) Eliminate J and solve for M

OpenStudy (anonymous):

Nice work, thank you all!!

OpenStudy (aum):

You are welcome.

OpenStudy (wolf1728):

COIN PROBLEM t = twenty cent pieces n = 5 cent pieces (nickels) t + n = 104 .2 t + .05 n = $10.00 32 20¢ pieces 72 5¢ pieces 32*.2 = 6.40 72*.05 = 3.60 Total = 10.00

OpenStudy (anonymous):

Thank you Wolf1728

OpenStudy (wolf1728):

u r welcome :-)

OpenStudy (yanasidlinskiy):

By the way, @remo \(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!}\)

OpenStudy (anonymous):

john is twice as old as mary. 5 years ago he was 3 times as old. what is mary's age now? could anyone please show me how to solve this.